# Unleashing the Potential of Physics-Inspired Neural Message Passing: From GNNs to Transformers

### v1

@@ Abstract
Learning representations for structured data with certain geometries (observed or unobserved) is a fundamental challenge, wherein message passing neural networks (MPNNs) have become a de facto class of model solutions. In this paper, we propose an energy-constrained diffusion model as a principled interpretable framework for understanding the mechanism of MPNNs and navigating novel architectural designs. The model, inspired by physical systems, combines the inductive bias of diffusion on manifolds with layer-wise constraints of energy minimization. As shown by our analysis, the diffusion operators have a one-to-one correspondence with the energy functions implicitly descended by the diffusion process, and the finite-difference iteration for solving the energy-constrained diffusion system induces the propagation layers of various types of MPNNs operated on observed or latent structures. On top of these findings, we devise a new class of neural message passing models, dubbed as diffusion-inspired Transformers, whose global attention layers are induced by the principled energy-constrained diffusion. Across diverse datasets ranging from real-world networks to images and physical particles, we show that the new model can yield promising performance for cases where the data structures are observed (as a graph), partially observed or completely unobserved.
@@ 1.1.1 Neural Diffusion on Graphs
Another line of research investigates PDE-inspired learning using the diffusion perspective as a principled guideline on top of which (discrete) neural network-based approaches are designed for node classification (Atwood and Towsley, 2016; Klicpera et al., 2019b; Xu et al., 2020), addressing over-smoothing (Rusch et al., 2023), knowledge distillation (Yang et al., 2022) and topological generalization (Wu et al., 2023). Our work leans on PDE-inspired learning and introduces a new diffusion model that is implicitly defined as minimizing a regularized energy. Our analysis also reveals the underlying equivalence between the numerical iterations of diffusion equations and unfolding the minimization dynamics of a corresponding energy. The results illuminate the fundamental connection between graph diffusion equations and energy optimization systems. More importantly, as we will show in the following sections, such a principled perspective (from the energy-constrained diffusion) brings up an interpretable framework for existing message passing neural networks and can be used for navigating new architecture designs.
@@ 1.1.2 Message Passing Neural Networks
Furthermore, from an architectural view, the majority of existing MPNNs operates the message passing per layer within observed edges of input graphs, which could limit efficacy in scenarios where the graphs are noisy or incomplete. To resolve this, several recent works propose to learn latent graph structures from data that can boost MPNNs towards better representations (Franceschi et al., 2019; Chen et al., 2020c; Jiang et al., 2019; Fatemi et al., 2021; Wu et al., 2022). These approaches generalize message-passing-based schemes to broader regimes where interactions are modeled by latent structures. Our work aims to provide a theoretical framework that can interpret the message passing rules of GNNs as numerical iterations of a diffusion process that descends a regularized energy in an interpretable form. As we will show in later sections, this principled perspective can be utilized to help understand the behavior of various MPNNs and the mechanism of message passing over observed or latent graph structures.
@@ 2 Preliminary and Background
where $F(u,t)$ denotes the diffusivity, $\odot$ denotes the Hadamard product, and Eqn. 1 can be incorporated with additional boundary conditions if $\Omega$ has a boundary. The physical implication of Eqn. 1 is that the temporal change of $z(u,t)$ at location $u$ equals to the flow that spatially enters into $u$ within infinitesimal time. The diffusivity determines the evolutionary direction of the diffusion system. If the diffusion is homogeneous, $F(u,t)$ remains a constant scalar for arbitrary $u$ and $t$. For non-homogeneous systems, the diffusion can be isotropic ($F(u,t)$ becomes a scalar-valued function and is location-dependent) and anisotropic ($F(u,t)$ becomes a $n\times n$ matrix-valued function and is location- and direction-dependent) (Weickert et al., 1998).
Another concept intimately related to diffusion is the energy that is a measure of how variable the physical quantity is in the system (Evans, 1998). For the quantity $z(u)$ on $\Omega$, the Dirichlet energy is a quadratic functional on the Sobolev space and returns a real number
@@ 3 Model Formulation: Energy-Constrained Diffusion
In this section, we introduce the general formulation of our energy-constrained diffusion model and its inherent connection with neural message passing. We will begin with a geometric diffusion model which is characterized by a diffusion PDE equation with flexible instantiations of the diffusivity. The latter enables us to bridge the numerical iterations of the PDE with different types of MPNNs. Built upon this, we will probe how the energy minimization perspective can be organically incorporated as a physics-inspired prior of the diffusion system in the form of constraints, which gives rise to the energy-constrained diffusion.
@@ 3.1 Geometric Diffusion with Observed/Latent Structures
The starting point of our model is rooted on an analogy that treats nodes in the graph (i.e., $i\in\mathcal{V}$) as locations on a Riemannian manifold (i.e., $u\in\Omega$(Rosenberg and Steven, 1997), node embeddings as the physical quantity of interest and the update of node embeddings per layer as heat flux through time (Chamberlain et al., 2021a).
@@ 5.2.1 Attention Designs Inspired by Time-Dependent Diffusivity
We next go into model instantiations based on the above theory, and introduce two specified $f$’s as practical versions of our model. To begin with, because $\|\mathbf{z}_{i}-\mathbf{z}_{j}\|_{2}^{2}=\|\mathbf{z}_{i}\|_{2}^{2}+\|\mathbf{z}_{j}\|_{2}^{2}-2\mathbf{z}_{i}^{\top}\mathbf{z}_{j}$, we can convert $f(\|\mathbf{z}_{i}-\mathbf{z}_{j}\|_{2}^{2})$ into the form $g(\mathbf{z}_{i}^{\top}\mathbf{z}_{j})$ on the condition that $\|\mathbf{z}_{i}\|_{2}$ remains constant. Notice that this condition is relatively mild since the normalization is often used to rescale the node embeddings before attention in practice. In particular, we use the L2 normalization to constrain $\mathbf{z}_{i}$ to have unit norm. Simple Diffusivity Attention. A straightforward design is to adopt the linear function $g(x)=1+x$ that gives rise to the dot-product attention:
@@ 5.3.2 Incorporating Graph Inductive Bias
The model presented in Section 5.2.1 does not assume any observed graph as input. For situations with observed structures (e.g., in the form of graphs), we can leverage the structural information as a geometric prior. Denote by $\mathcal{G}=(\mathcal{V},\mathcal{E})$ the input graph, and here we adopt a simple-yet-effective design by modifying the layer-wise updating rule as:
@@ 5.3.4 Scaling Up DIFFormer to Large-Scale Systems
One remaining issue for our model is how to scale the global attention to large-scale systems that involve structures (observed or unobserved) among massive numbers of nodes, e.g., up to millions. The large number of inter-connected points makes it hard for full-batch training on a single GPU. However, thanks to the reduced reliance on input graphs, we can harness a simple strategy for improving the space efficiency for training DIFFormer. To be specific, in each epoch, we partition the dataset into random mini-batches and feed one mini-batch (including the input features of nodes within the current mini-batch and if any, the graph adjacency composed by the observed structures among these nodes) for one feed-forward and backward computation. In particular, for DIFFormer-s that only requires linear complexity w.r.t. node numbers, we can set a large batch size in practice which gives rise to enough global information of the interactions among nodes in one mini-batch. The flexibility of mini-batch training also accommodates parallel acceleration if needed in practice.
@@ 6.1 Learning with Observed Structures
Results on Homophilous Graphs. We test DIFFormer on three citation networks Cora, Citeseer and Pubmed which are commonly used as benchmarks for evaluating graph representation learning approaches. Table 2 reports the testing accuracy. We compare with several sets of baselines linked with our model from different aspects. The first category of competitors includes basic models: MLP (Rumelhart et al., 1986) and two classical graph-based semi-supervised learning approaches Label Propagation (LP(Zhu et al., 2003) and ManiReg (Belkin et al., 2006). The second category of competitors belongs to MPNNs on observed graphs, including commonly used GNNs (SGC (Wu et al., 2019), GCN (Kipf and Welling, 2017) and GAT (Velickovic et al., 2018)), diffusion-based models (GRAND (Chamberlain et al., 2021a), its linear variant GRAND-l, and GRAND++ (Thorpe et al., 2022)), and diffusion-inspired models (GDC (Klicpera et al., 2019b) and GraphHeat (Xu et al., 2020). The third model category extends MPNNs to latent graphs. In particular, we consider MPNNs operated on $k$-nearest-neighbor graphs (GCN-$k$NN and GAT-$k$NN) and latent complete graphs that connect all node pairs (Dense GAT). Furthermore, we compare with several strong structure learning models that learn to optimize the latent structures on which MPNNs run, including LDS (Franceschi et al., 2019), GLCN (Jiang et al., 2019), and NodeFormer (Wu et al., 2022). Table 2 shows that DIFFormer achieves the best results on three datasets with significant improvements. Also, we notice that the simple diffusivity model DIFFormer-s significantly exceeds the counterparts without non-linearity (SGC, GRAND-l and DGC-Euler) and even comes to the first on Cora and Pubmed. These results suggest that DIFFormer can serve as a very competitive encoder backbone for node-level prediction that learns inter-instance interactions for generating informative representations and boosting downstream performance.
@@ 7 Conclusions
The authors would like to thank the anonymous reviewers of ICLR 2023 as well as Prof. Michael Bronstein, Prof. Hongyuan Zha and Prof. Shi Jin for their insightful comments and suggestions on this work.

### v2

@@ Abstract
Learning representations for structured data with certain geometries (e.g., observed or unobserved) is a fundamental challenge, wherein message passing neural networks (MPNNs) have become a de facto class of model solutions. In this paper, we propose an energy-constrained diffusion model as a principled mathematical framework for understanding the mechanism of MPNNs and navigating novel architectural designs. Inspired by physical systems, the model combines the inductive bias of diffusion on manifolds with layer-wise constraints of energy minimization. We identify that the diffusion operators have a one-to-one correspondence with the energy functions implicitly descended by the diffusion process, and the finite-difference iteration for solving the energy-constrained diffusion system induces the propagation layers of various types of MPNNs operating on observed or latent structures. This leads to a unified perspective on common neural architectures whose computational flows can be cast as message passing (or its special case), including MLP, GCN, GIN, APPNP, GCNII, GAT, and Transformers. Building on these insights, we devise a new class of neural message passing models, dubbed diffusion-inspired Transformers, whose global attention layers are derived from the principled energy-constrained diffusion framework. Across diverse datasets—from real-world networks to images, texts, and physical particles—we demonstrate that the new model achieves promising performance in scenarios where the data structures are observed (as a graph), partially observed, or entirely unobserved. [^1] Code available at https://github.com/qitianwu/DIFFormer
@@ 1.1.1 Neural Diffusion on Graphs
Another line of research investigates PDE-inspired learning using the diffusion perspective as a principled guideline on top of which (discrete) neural network-based approaches are designed for node classification (Atwood and Towsley, 2016; Klicpera et al., 2019b; Xu et al., 2020), addressing over-smoothing (Rusch et al., 2023), knowledge distillation (Yang et al., 2022) and topological generalization (Wu et al., 2025). Our work leans on PDE-inspired learning and introduces a new diffusion model that is implicitly defined as minimizing a regularized energy. Our analysis also reveals the underlying equivalence between the numerical iterations of diffusion equations and unfolding the minimization dynamics of a corresponding energy. The results illuminate the fundamental connection between graph diffusion equations and energy optimization systems. More importantly, as we will show in the following sections, such a principled perspective (from the energy-constrained diffusion) brings up an interpretable framework for existing message passing neural networks and can be used for navigating new architecture designs.
@@ 1.1.2 Message Passing Neural Networks
Furthermore, from an architectural view, the majority of existing MPNNs operates the message passing per layer within observed edges of input graphs, which could limit efficacy in scenarios where the graphs are noisy or incomplete. To resolve this, several recent works propose to learn latent graph structures from data that can boost MPNNs towards better representations (Franceschi et al., 2019; Chen et al., 2020c; Jiang et al., 2019; Fatemi et al., 2021; Wu et al., 2022, 2023a, 2023b; Deng et al., 2024). These approaches generalize message-passing-based schemes to broader regimes where interactions are modeled by latent structures. Our work aims to provide a theoretical framework that can interpret the message passing rules of GNNs as numerical iterations of a diffusion process that descends a regularized energy in an interpretable form. As we will show in later sections, this principled perspective can be utilized to help understand the behavior of various MPNNs and the mechanism of message passing over observed or latent graph structures.
@@ 1.2 Contributions and Organization
Comparison with the Conference Paper on ICLR 2023 (Wu et al., 2023a). On the basis of our conference paper, we have made substantial extensions that entail new theoretical analysis and empirical results enriching and deepening the technical contents. In Section 2, we add more technical background about the manifold diffusion as foundations of our proposed model. In Section 4.1, we analyze linear diffusion equations with constant diffusivity, and show its connection with the minimization of the convex energy function of quadratic forms. We also illustrate how to derive various types of GNNs, such as GCN, GIN and APPNP, starting from the energy-constrained diffusion with constant diffusivity. In Section 4.2, we derive the upper and lower bound of the energy at each layer and discuss how to avoid the over-smoothing issue with a source term incorporated into the diffusion equation. In Section 5.2.2, we present discussions on how to derive dot-then-exponential Softmax attention via the energy-constrained diffusion with time-dependent diffusivity. And in Section 5.3, we supplement detailed discussions along with analysis on how to extend our model with feature transformations, non-linear activations, and graph inductive biases. For the experiments, we add empirical comparisons on five additional datasets including heterophilic graphs (in Section 6.1) and physical particles (in Section 6.3) and discussions on addressing the potential over-smoothing (in Section 6.4).
@@ 2 Preliminary and Background
where $F(u,t)$ denotes the diffusivity, $\odot$ denotes the Hadamard product, and Eqn. 1 can be incorporated with additional boundary conditions if $\Omega$ has a boundary. The physical implication of Eqn. 1 is that the temporal change of $z(u,t)$ at location $u$ equals to the flow that spatially enters into $u$ within infinitesimal time. This resembles the main design of message passing neural networks (MPNNs), where as layers increase (time goes by), the embeddings (physical quantity) of connected nodes (adjacent locations) are propagated to update that of each other. More illustration on how these two perspectives are bridged through analogy is outlined in Fig. 2 and will be elaborated in Sec. 3.
The diffusivity in diffusion equations determines the evolutionary direction of the diffusion system. If the diffusion is homogeneous, $F(u,t)$ remains a constant scalar for arbitrary $u$ and $t$. For non-homogeneous systems, the diffusion can be isotropic ($F(u,t)$ becomes a scalar-valued function and is location-dependent) and anisotropic ($F(u,t)$ becomes a $n\times n$ matrix-valued function and is location- and direction-dependent) (Weickert et al., 1998). As will be introduced in later sections, different instantiations of diffusivity will give rise to specific forms of MPNNs such as MLP, GCN, GAT and Transformers.
Another concept intimately related to diffusion is the energy, a measure of how variable the physical quantity is in the system (Evans, 1998). For the quantity $z(u)$ on $\Omega$, the Dirichlet energy is a quadratic functional on the Sobolev space and returns a real number
@@
Figure 2: Illustration of (a) the connection between diffusion process on manifolds and message passing neural networks (MPNNs), (b) the analogy bridging the concepts of two models, and (c) definitions of key components in diffusion equation models on graphs.
@@ 3 Model Formulation: Energy-Constrained Diffusion
In this section, we introduce the general formulation of our energy-constrained diffusion model and its inherent connection with neural message passing. We will begin with a geometric diffusion model which is characterized by a diffusion PDE equation with flexible instantiations of the diffusivity. The latter enables us to bridge the numerical iterations of the PDE with different types of MPNNs. Built upon this, we will probe how the energy minimization perspective can be organically incorporated as a physics-inspired prior of the diffusion system in the form of constraints, which gives rise to energy-constrained diffusion.
@@ 3.1 Geometric Diffusion with Observed/Latent Structures
The starting point of our model is rooted on an analogy that treats nodes in the graph (i.e., $i\in\mathcal{V}$) as locations on a Riemannian manifold (i.e., $u\in\Omega$(Rosenberg and Steven, 1997), node embeddings as the physical quantity of interest and the update of node embeddings per layer as heat flux through time (Chamberlain et al., 2021a). A high-level illustration is presented in Fig. 2.
@@ 5.2.1 Attention Designs Inspired by Time-Dependent Diffusivity
We next go into model instantiations based on the above theory, and introduce two specified $f$’s as practical versions of our model. To begin with, because $\|\mathbf{z}_{i}-\mathbf{z}_{j}\|_{2}^{2}=\|\mathbf{z}_{i}\|_{2}^{2}+\|\mathbf{z}_{j}\|_{2}^{2}-2\mathbf{z}_{i}^{\top}\mathbf{z}_{j}$, we can convert $f(\|\mathbf{z}_{i}-\mathbf{z}_{j}\|_{2}^{2})$ into the form $g(\mathbf{z}_{i}^{\top}\mathbf{z}_{j})$ on the condition that $\|\mathbf{z}_{i}\|_{2}$ remains constant. Notice that this condition is relatively mild since the normalization is often used to rescale the node embeddings before attention in practice. In particular, we use the L2 normalization to constrain $\mathbf{z}_{i}$ to have unit norm.
Simple Diffusivity Attention. A straightforward design is to adopt the linear function $g(x)=1+x$ that gives rise to the dot-product attention:
@@ 5.3.2 Incorporating Graph Inductive Bias
The model presented in Section 5.2.1 does not assume any observed graph as input. For situations with observed structures (e.g., in the form of graphs), we can leverage the structural information as a geometric prior. There potentially exist different ways to utilize the graph information. For instance, Dwivedi and Bresson (2020) uses the features of Laplacian decomposition on the input graphs as absolute positional embeddings, and incorporates the positional embeddings with node features as the input for Transformers. This scheme, however, is not scalable for large graphs due to the $\mathcal{O}(|\mathcal{V}|^{3})$ complexity of Laplacian decomposition. Furthermore, Ying et al. (2021) considers the encodings of node-pair-wise distances as relative positional embeddings that are used to reinforce the attention weight for any node pair. Since the distance encodings introduce extra trainable parameters and involve arbitrary node pairs in the graph (whose numbers scale quadratically w.r.t. the graph size), this approach significantly increases the training cost on large graphs. In consideration of both effectiveness and efficiency criteria, we turn to a simple-yet-effective scheme for incorporating the graph structural information. Denote by $\mathcal{G}=(\mathcal{V},\mathcal{E})$ the input graph, and we modify the layer-wise updating rule as:
@@ 5.3.4 Scaling Up DIFFormer to Large-Scale Systems
One remaining issue for our model is how to scale the global attention to large-scale systems that involve structures (observed or unobserved) among massive numbers of nodes, e.g., up to millions. The large number of inter-connected points makes it hard for full-batch training on a single GPU. However, thanks to the reduced reliance on input graphs, we can harness a simple strategy for improving the space efficiency for training DIFFormer. To be specific, in each epoch, we partition the dataset into random mini-batches and feed one mini-batch (including the input features of nodes within the current mini-batch and if any, the graph adjacency composed by the observed structures among these nodes) for one feed-forward and backward computation. In particular, for DIFFormer-s that only requires linear complexity w.r.t. node numbers, we can set a large batch size in practice which gives rise to enough global information of the interactions among nodes in one mini-batch. Also, for common large graphs with millions of nodes, the mini-batch partition is only required for training, so the global interactions can be fully accommodated at test time. The flexibility of mini-batch training also accommodates parallel acceleration if needed in practice.
@@ 5.4 Systematic Comparisons of Different Model Classes
Apart from these, the proposed model framework is also related to Label Propagation (Zhou et al., 2003) which is a classic semi-supervised learning algorithm that propagates the labels of training samples to predict those of testing samples. One straightforward way to bridge both of the worlds is to replace the node embeddings $\mathbf{Z}$ with labels $\mathbf{Y}$, on top of which one can derive the diffusion process and energy functions induced by different label propagation algorithms. Pushing further, a recent work (Yang et al., 2024) identifies the inherent connection between MPNNs and label propagation, where the training dynamics of MPNNs can be written as a form of label propagation and the propagation matrix evolves during training. This work sheds lights on another perspective that dissects the optimization process of MPNNs. Though in general the results of Yang et al. (2024) are orthogonal to our work, our analysis can be extended to investigate into the optimization dynamics of different MPNN models during training and reveal their implicit bias by linking the label propagation view with energy-constrained diffusion.
@@ 6.1 Learning with Observed Structures
Results on Homophilous Graphs. We test DIFFormer on three citation networks Cora, Citeseer and Pubmed which are commonly used as benchmarks for evaluating graph representation learning approaches. Table 2 reports the testing accuracy. We compare with several sets of baselines linked with our model from different aspects. The first category of competitors includes basic models: MLP (Rumelhart et al., 1986) and two classical graph-based semi-supervised learning approaches Label Propagation (LP(Zhu et al., 2003) and ManiReg (Belkin et al., 2006). The second category of competitors belongs to MPNNs on observed graphs, including commonly used GNNs (SGC (Wu et al., 2019), GCN (Kipf and Welling, 2017) and GAT (Velickovic et al., 2018)), diffusion-based models (GRAND (Chamberlain et al., 2021a), its linear variant GRAND-l, and GRAND++ (Thorpe et al., 2022)), and diffusion-inspired models (GDC (Klicpera et al., 2019b) and GraphHeat (Xu et al., 2020). The third model category extends MPNNs to latent graphs. In particular, we consider MPNNs operated on $k$-nearest-neighbor graphs (GCN-$k$NN and GAT-$k$NN) and latent complete graphs that connect all node pairs (Dense GAT). Furthermore, we compare with several strong structure learning models that learn to optimize the latent structures on which MPNNs run, including LDS (Franceschi et al., 2019), GLCN (Jiang et al., 2019), and three recently proposed graph Transformer models including NodeFormer (Wu et al., 2022), Graphormer (Ying et al., 2021) and GraphGPS (Rampášek et al., 2022). Table 2 shows that DIFFormer achieves the best results on three datasets with significant improvements. Also, we notice that the simple diffusivity model DIFFormer-s significantly exceeds the counterparts without non-linearity (SGC, GRAND-l and DGC-Euler) and even comes to the first on Cora and Pubmed. These results suggest that DIFFormer can serve as a very competitive encoder backbone for node-level prediction that learns inter-instance interactions for generating informative representations and boosting downstream performance.
@@ 7 Conclusions
Qitian Wu and Junchi Yan were partly supported by NSFC (92370201, 62222607). The authors would like to thank the anonymous reviewers of ICLR 2023 as well as Michael Bronstein, Hongyuan Zha and Shi Jin for their insightful comments and suggestions on this work.